Aarushi has a total of RS 590 as currency notes in the denominations of RS 50, RS 20 and RS 10. The ratio of number of RS 50 and RS 20 is 3:5. If she has total of 25 notes, how many notes of each denomination she has?

Answer by stanbon(75887)

Aarushi has a total of RS 590 as currency notes in the denominations of RS 50, RS 20 and RS 10. The ratio of number of RS 50 and RS 20 is 3:5. If she has total of 25 notes, how many notes of each denomination she has?
-------
Value:: 50f + 20t + 10n = 590
Quantity:: 5f = 3t
Quantity:: f + t + n = 25
-------------------------------------
Modify to solve::
f = 3/5 t
-----
Substitute to get:
30t + 20t + 10n = 590
(8/5)t + n = 25
---------------------
50t + 10n = 590
8t + 5n = 125
--------------------
50t + 10n = 590
16t + 10n = 250
-------------------
Subtract and solve for "t":
34t = 340
t = 10 (# of RS20 bills)
---
f = (3/5)t = 6 (# of RS 50 bills)
-----
n = 25-(8/5)t = 25-16 = 9 (# of RS10 bills)
-------------
Cheers,
Stan H.
-------------